Calculate the number of charge of a metal which passes through a solution with 2.3A for 30minutes oand the mass of the metal is given to be 3.6g.

Molar m. =65gmol-¹
1
Expert's answer
2020-01-08T02:02:08-0500

Q=I×t=2.3×(30×60)=4140 (C)

Answer: 4140 C.